# Classification of compact manifolds with positive isotropic curvature

### v6

@@ 2 $\varepsilon$-caps and ancient $\kappa$-solutions
Let $\mathcal{O}^{\prime}$ be a compact suborbifold of an orbifold $\mathcal{O}$, and $\varphi:D(\xi)\rightarrow\mathcal{O}$ and $\psi:D(\eta)\rightarrow\mathcal{O}$ be two closed tubular neighborhoods of $\mathcal{O}^{\prime}$. Then there exists an isotopy $H_{t}$, $t\in[0,1]$, of $\mathcal{O}$ fixing $\mathcal{O}^{\prime}$ pointwisely, with compact support, and with $H_{0}=\text{id}$ and $H_{1}\circ\varphi=\psi\circ\Lambda$, where $\Lambda:D(\xi)\rightarrow D(\eta)$ is an orthogonal bundle isomorphism.
Proof.   First note that the two closed tubular neighborhoods $\varphi$ and $\psi$ are isotopic. This can be shown by using local models for the orbifold vector bundles $\xi$ and $\eta$; cf. the proof of Theorem 2.6 in Chapter VI in [6] and Theorem 7.4.6 in [29]. Alternatively, as pointed out to me by Prof. Bonahon, one can also show this by using normal exponential maps and the fact that the space of Riemannian metrics on the orbifold $\mathcal{O}$ is convex; cf. Section 6.1 of [2]. Then we can apply a version of Thom’s isotopy extension theorem for orbifolds. The (statement and) proof of the orbifold version of Thom’s isotopy extension theorem can be adapted from that of Theorem 9.5 in [8] and Theorem 7.3.3 in [29] (cf. also p. 179-181 of [25] and Theorem 2.4.2 in [37]). More precisely, the desired ambient isotopy in the orbifold version of Thom’s isotopy extension theorem comes from the projection of a suitable global flow on ×$\mathbb{R}\times\mathcal{O}$, the latter being generated by a vector field on ×$\mathbb{R}\times\mathcal{O}$. To construct this vector field, we first construct suitable vector fields in local models of ×$\mathbb{R}\times\mathcal{O}$ as in the proof of Theorem 9.5 in [8], Theorem 7.3.3 in [29] and Theorem 2.4.2 in [37], then get invariant vector fields in local models by using the method of averaging as in the proof of Theorem 3.1 in Chapter VI in [6], finally glue these local vector fields to a global vector field via a partition of unity. $\Box$
@@
Let $\Gamma$ be as above, and $\mathcal{O}$ be a $n$-dimensional orbifold diffeomorphic to $\mathbb{S}^{n}//\Gamma$. Write $\mathcal{O}=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{2}$ is diffeomorphic to $D_{-}$, and $\Omega_{1}\cap\Omega_{2}=\partial\Omega_{1}=\partial\Omega_{2}$. Assume that $h:\partial D_{-}\rightarrow\partial\Omega_{2}$ is a diffeomorphism which extends to a diffeomorphism $\tilde{h}:D_{-}\rightarrow\Omega_{2}$. Then there is a $\lambda\in\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ such that $h\circ\lambda$ extends to a diffeomorphism $\tilde{h}^{\prime}:D_{+}\rightarrow\Omega_{1}$.
Proof.   Let $f:\mathbb{S}^{n}//\Gamma\rightarrow\mathcal{O}$ be a diffeomorphism. The case that $\Gamma$ is trivial follows from the Cerf-Palais disk theorem, and is well-known. Now we assume that $\Gamma$ is non-trivial. W.l.o.g. we may assume that $f(p_{-})\in\Omega_{2}$, since otherwise we can replace $f$ by $f$ precomposing with an isometric involution of $\mathbb{S}^{n}//\Gamma$ which exchanges $D_{+}$ and $D_{-}$. Then we may view both $f|_{D_{-}}:D_{-}\rightarrow f(D_{-})$ and $\tilde{h}:D_{-}\rightarrow\Omega_{2}$ as a closed tubular neighborhood of the singular point $f(p_{-})=\tilde{h}(p_{-})$. By the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds (Proposition 2.1) there exist a $\Lambda\in\text{Isom}(D_{-})$ and a self diffeomorphism $F$ of $\mathcal{O}$ such that $F\circ f|_{D_{-}}=\tilde{h}\circ\Lambda$. Then $F\circ f|_{D_{+}}:D_{+}\rightarrow\Omega_{1}$ extends $h\circ\Lambda|_{\mathbb{S}^{n-1}/\Gamma}$, and we are done.
$\Box$
@@
Let $\Gamma$, $\sigma$, and $\hat{\sigma}$ be as before, and $\mathcal{O}$ be a $n$-dimensional orbifold which is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$. Write $\mathcal{O}=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{2}$ is diffeomorphic to $\bar{B}$, and $\Omega_{1}\cap\Omega_{2}=\partial\Omega_{1}=\partial\Omega_{2}$. Assume that $h:\partial\bar{B}\rightarrow\partial\Omega_{2}$ is a diffeomorphism which extends to a diffeomorphism $\tilde{h}:\bar{B}\rightarrow\Omega_{2}$. Then there is a $\lambda\in\text{Isom}(\partial\bar{B})$ such that $h\circ\lambda$ extends to a diffeomorphism $\tilde{h}^{\prime}:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B\rightarrow\Omega_{1}$.
@@
Now we start to analyze the structure of orbifold ancient $\kappa$-solutions which satisfy certain pinching assumption. (Recall from [7] and [19] that the orbifold ancient $\kappa$-solutions we consider are weakly PIC2.) Let $r_{1}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be the reflection w.r.t. the hyperplane $x_{1}=0$, we also use $r_{1}$ to denote the restriction of $r_{1}$ to $\mathbb{S}^{n-1}$. The following result is an improvement of Proposition 2.5 in [19]. (Recall that $C_{\text{PIC}}$, $C_{\text{PIC}1}$ and $C_{\text{PIC}2}$ are the sets of algebraic curvature tensors which have nonnegative isotropic curvature (weakly PIC, see item (i) of Definition 1.1 in [7]), weakly PIC1 (see item (ii) of Definition 1.1 in [7]), and weakly PIC2 (see item (iii) of Definition 1.1 in [7]) respectively.)
@@
(cf. Corollary 6.7 in [7], Theorem 3.4 in [10], and Proposition 2.5 in [19])    Let $n\geq 4$, and $(\mathcal{O},g(t))$, $t\in(-\infty,T]$, be an orbifold ancient $\kappa$-solution of dimension $n$ with at most isolated singularities. Suppose that $(\mathcal{O},g(t))$ satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ for some uniform constant $\theta>0$, and there is a spacetime point $(x_{0},t_{0})$ such that the curvature tensor at $(x_{0},t_{0})$ lies on the boundary of the cone $C_{\text{PIC}2}$. Then for each $t$, $(\mathcal{O},g(t))$ is isometric to a shrinking Ricci soliton $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ or $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$. In particular, if $\mathcal{O}$ has no singularities and has exactly one end, it must be diffeomorphic to some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with no fixed points, and $\hat{\sigma}$ and $B$ are defined as before; if $\mathcal{O}$ has nonempty isolated singularities, it must be diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$ defined before. Furthermore, there exists a positive constant $\varepsilon_{4}=\varepsilon_{4}(n)\leq\varepsilon_{1}$, where $\varepsilon_{1}$ is the constant in Proposition 5.4 in [19], with the following property.
@@
As in the proof of Proposition 2.5 in [19], we can write $\mathcal{O}=\mathbb{S}^{n-1}\times\mathbb{R}/\widetilde{\Gamma}$, where $\widetilde{\Gamma}$ is a subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ with $\widetilde{\Gamma}=\Gamma\cup\Gamma^{1}$, where the second component of $\Gamma$ (resp. of $\Gamma^{1}$) acts on $\mathbb{R}$ as the identity (resp. a reflection). Pick $\tilde{\sigma}\in\Gamma^{1}$. Then $\tilde{\sigma}^{2}\in\Gamma$, and $\tilde{\sigma}\Gamma=\Gamma^{1}$. So $\tilde{\sigma}$ induces an isometric involution on $\mathbb{S}^{n-1}/\Gamma$ (without any fixed points by our assumption on $\mathcal{O}$), which will be denoted by $\sigma$. The action of $\tilde{\Gamma}$ on $\mathbb{S}^{n-1}\times\mathbb{R}$ will leave exactly one cross section, say $\mathbb{S}^{n-1}\times\{0\}$, invariant. (Compare the beginning of the third paragraph on p. 53 in [10].) Let $\pi:\mathbb{S}^{n-1}\times\mathbb{R}\rightarrow\mathcal{O}=\mathbb{S}^{n-1}\times\mathbb{R}/\tilde{\Gamma}$ be the natural projection. Then $\pi(\mathbb{S}^{n-1}\times\{0\})$ is a homothetically embedded $(\mathbb{S}^{n-1}/\Gamma)/\langle\sigma\rangle$ in $(\mathcal{O},g(t))$. Fix any $t$ and write $g=g(t)$. From Proposition 5.4 in [19] we have the following
@@
(cf. Proposition 6.17 in [7] and Proposition 2.6 in [19]) Let $n\geq 4$ and $\varepsilon_{4}=\varepsilon_{4}(n)$ be as in Proposition 2.4. Let $(\mathcal{O},g)$ be a complete, noncompact orbifold diffeomorphic to $\mathbb{R}^{n}//\Gamma$ with strictly PIC and weakly PIC2 everywhere and strictly PIC2 at a point. Assume that $N$ is a $2\varepsilon_{4}$-neck in $\mathcal{O}$ diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$, and $\Sigma$ is a central cross section of $N$. When $\Gamma$ is nontrivial we assume that $\Sigma$ bounds a compact domain $\Omega$ in $\mathcal{O}$ whose interior contains the unique singular point of $\mathcal{O}$. Then if $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ is a $2\varepsilon_{4}$-homothety coming from the $2\varepsilon_{4}$-neck structure of $N$, there exists a $\lambda\in\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ such that $f\circ\lambda$ extends to a diffeomorphism $F:D^{n}//\Gamma\rightarrow\Omega$.
@@
Definition. (cf. [7] and [19]) Let $\varepsilon_{0}=\varepsilon_{0}(n)$ be a small positive constant and $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$. Let $(\mathcal{O},g)$ be a $n$-dimensional Riemannian orbifold with at most isolated singularities. Given a point $x_{0}\in\mathcal{O}$, an open subset $U$ of $\mathcal{O}$ is an $\varepsilon$-cap centered at $x_{0}$ if $U$ is a topological cap and $U\setminus V$ is an $\varepsilon$-neck, where $V$ is a compact domain in $U$ diffeomorphic to some $D^{n}//\Gamma$ or $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B$ or $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B$ with $\partial V$ a central cross section of an $\varepsilon$-neck contained in $U$ (i.e., $\partial V$ is the image of $\mathbb{S}^{n-1}/\Gamma\times\{0\}$ under the diffeomorphism associated to the $\varepsilon$-neck) and with $x_{0}\in\text{Int}\hskip 1.42262ptV$, (we will call $V$ a core of the cap $U$,) and in addition,
1. when $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$, if $\Sigma$ is a central cross section of an $\varepsilon_{0}$-neck in $U\setminus V$ with an $\varepsilon_{0}$-homothety $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ coming from the $\varepsilon_{0}$-neck structure, $\Sigma$ bounds a compact domain $\Omega$ in $U$, and $f\circ\lambda$ extends to a diffeomorphism $F:D^{n}//\Gamma\rightarrow\Omega$ for some isometry $\lambda$ of $\mathbb{S}^{n-1}/\Gamma$;
@@
(cf. Theorem 6.18 in [7], Theorem 3.9 in [10], and Proposition 2.7 in [19])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$ and $C_{2}=C_{2}(n,\theta,\varepsilon)$, such that given any noncompact orbifold ancient $\kappa$-solution $(\mathcal{O},g(t))$ of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ and is not locally isometric to an evolving shrinking round cylinder, for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$, which falls into one of the following categories:
(a) $U$ is an $\varepsilon$-neck centered at $(x_{0},t_{0})$,
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ and diffeomorphic to $\mathbb{R}^{n}//\Gamma$;
moreover, the scalar curvature in $U$ at time $t_{0}$ is (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$.

### v8

@@ 2 $\varepsilon$-caps and ancient $\kappa$-solutions
Let $\mathcal{O}^{\prime}$ be a compact embedded full suborbifold (without boundary, not necessarily connected) of an orbifold $\mathcal{O}$, and $\varphi:D(\xi)\rightarrow\mathcal{O}$ and $\psi:D(\eta)\rightarrow\mathcal{O}$ be two closed tubular neighborhoods of $\mathcal{O}^{\prime}$. Then there exists an isotopy $H_{t}$, $t\in[0,1]$, of $\mathcal{O}$ fixing $\mathcal{O}^{\prime}$ pointwisely, with compact support, and with $H_{0}=\text{id}$ and $H_{1}\circ\varphi=\psi\circ\Lambda$, where $\Lambda:D(\xi)\rightarrow D(\eta)$ is an orthogonal bundle isomorphism.
Proof. First note that the two closed tubular neighborhoods $\varphi$ and $\psi$ are isotopic. This can be shown by using local models for the orbifold vector bundles $\xi$ and $\eta$; cf. the proof of Theorem 2.6 in Chapter VI in [10], Theorem 4.6 in [28], and Theorem 7.4.6 in [39]. Alternatively, as pointed out to me by Prof. Bonahon, one can also show this by using normal exponential maps and the fact that the space of Riemannian metrics on the orbifold $\mathcal{O}$ is convex; cf. Section 6.1 of [3]. Then we can apply Lemma 2.1 (in the case that the embedded full suborbifold $\mathcal{P}$ has codimension 0 and has boundary). $\Box$
@@
Let $\Gamma$ be as above, and $\mathcal{O}$ be an $n$-dimensional orbifold diffeomorphic to $\mathbb{S}^{n}//\Gamma$. Write $\mathcal{O}=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{2}$ is diffeomorphic to $D_{-}$, and $\Omega_{1}\cap\Omega_{2}=\partial\Omega_{1}=\partial\Omega_{2}$. Assume that $h:\partial D_{-}\rightarrow\partial\Omega_{2}$ is a diffeomorphism which extends to a diffeomorphism $\tilde{h}:D_{-}\rightarrow\Omega_{2}$. Then $h$ extends to a diffeomorphism $\tilde{h}^{\prime}:D_{+}\rightarrow\Omega_{1}$.
Proof. Let $f:\mathbb{S}^{n}//\Gamma\rightarrow\mathcal{O}$ be a diffeomorphism. The case that $\Gamma$ is trivial follows from the Cerf-Palais disk theorem, and is well-known. Now we assume that $\Gamma$ is non-trivial. W.l.o.g. we may assume that $f(p_{-})\in\Omega_{2}$, since otherwise we can replace $f$ by $f$ precomposing with an isometric involution of $\mathbb{S}^{n}//\Gamma$ which exchanges $D_{+}$ and $D_{-}$. Then we may view both $f|_{D_{-}}:D_{-}\rightarrow f(D_{-})$ and $\tilde{h}:D_{-}\rightarrow\Omega_{2}$ as a closed tubular neighborhood of the singular point $f(p_{-})=\tilde{h}(p_{-})$. By the ambient isotopy uniqueness of closed tubular neighborhoods of compact embedded full suborbifolds (Proposition 2.2) there exist a $\Lambda\in\text{Isom}(D_{-})$ and a self diffeomorphism $F$ of $\mathcal{O}$ such that $F\circ f|_{D_{-}}=\tilde{h}\circ\Lambda$. Then $F\circ f|_{D_{+}}:D_{+}\rightarrow\Omega_{1}$ extends $h\circ\Lambda|_{\mathbb{S}^{n-1}/\Gamma}$. But $(\Lambda|_{\mathbb{S}^{n-1}/\Gamma})^{-1}$ extends to an isometry of $D_{+}$, so we are done. $\Box$
@@
Let $\Gamma$, $\sigma$, and $\hat{\sigma}$ be as before, and $\mathcal{O}$ be an $n$-dimensional orbifold which is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$. Write $\mathcal{O}=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{2}$ is diffeomorphic to $\bar{B}$, and $\Omega_{1}\cap\Omega_{2}=\partial\Omega_{1}=\partial\Omega_{2}$. Assume that $h:\partial\bar{B}\rightarrow\partial\Omega_{2}$ is a diffeomorphism which extends to a diffeomorphism $\tilde{h}:\bar{B}\rightarrow\Omega_{2}$. Then there is a $\lambda\in\text{Isom}(\partial\bar{B})$ such that $h\circ\lambda$ extends to a diffeomorphism $\tilde{h}^{\prime}:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B\rightarrow\Omega_{1}$.
@@
Now we start to analyze the structure of orbifold ancient $\kappa$-solutions which satisfy a certain pinching assumption. (Recall from [11] and [27] that the orbifold ancient $\kappa$-solutions we consider are weakly PIC2.) Let $r_{1}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be the reflection w.r.t. the hyperplane $x_{1}=0$, we also use $r_{1}$ to denote the restriction of $r_{1}$ to $\mathbb{S}^{n-1}$. The following result is an improvement of Proposition 2.5 in [27]. (Recall that $C_{\text{PIC}}$, $C_{\text{PIC}1}$ and $C_{\text{PIC}2}$ are the sets of algebraic curvature tensors which have nonnegative isotropic curvature (weakly PIC, see item (i) of Definition 1.1 in [11]), weakly PIC1 (see item (ii) of Definition 1.1 in [11]), and weakly PIC2 (see item (iii) of Definition 1.1 in [11]) respectively.)
@@
(cf. Corollary 6.7 in [11], Theorem 3.4 in [15], and Proposition 2.5 in [27]) Let $n\geq 4$, and $(\mathcal{O},g(t))$, $t\in(-\infty,T]$, be an orbifold ancient $\kappa$-solution of dimension $n$ with at most isolated singularities. Suppose that $(\mathcal{O},g(t))$ satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}\text{id}\in C_{\text{PIC}}$ for some uniform constant $\theta>0$, and there is a spacetime point $(x_{0},t_{0})$ such that the curvature tensor at $(x_{0},t_{0})$ lies on the boundary of the cone $C_{\text{PIC}2}$. Then for each $t$, $(\mathcal{O},g(t))$ is isometric to a shrinking Ricci soliton $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ or $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$. In particular, if $\mathcal{O}$ has no singularities and has exactly one end, it must be diffeomorphic to some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with no fixed points, and $\hat{\sigma}$ and $B$ are defined as before; if $\mathcal{O}$ has nonempty isolated singularities, it must be diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$ defined before. Furthermore, there exists a positive constant $\varepsilon_{4}=\varepsilon_{4}(n)\leq\varepsilon_{1}$, where $\varepsilon_{1}$ is the constant in Proposition A.4 in [27], with the following property.
@@
As in the proof of Proposition 2.5 in [27], we can write $\mathcal{O}=\mathbb{S}^{n-1}\times\mathbb{R}/\widetilde{\Gamma}$, where $\widetilde{\Gamma}$ is a subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ with $\widetilde{\Gamma}=\Gamma\cup\Gamma^{1}$, where the second component of $\Gamma$ (resp. of $\Gamma^{1}$) acts on $\mathbb{R}$ as the identity (resp. a reflection). Pick $\tilde{\sigma}\in\Gamma^{1}$. Then $\tilde{\sigma}^{2}\in\Gamma$, and $\tilde{\sigma}\Gamma=\Gamma^{1}$. So $\tilde{\sigma}$ induces an isometric involution on $\mathbb{S}^{n-1}/\Gamma$ (without any fixed points by our assumption on $\mathcal{O}$), which will be denoted by $\sigma$. The action of $\tilde{\Gamma}$ on $\mathbb{S}^{n-1}\times\mathbb{R}$ will leave exactly one cross section, say $\mathbb{S}^{n-1}\times\{0\}$, invariant. (Compare the beginning of the third paragraph on p. 53 in [15].) Let $\pi:\mathbb{S}^{n-1}\times\mathbb{R}\rightarrow\mathcal{O}=\mathbb{S}^{n-1}\times\mathbb{R}/\tilde{\Gamma}$ be the natural projection. Then $\pi(\mathbb{S}^{n-1}\times\{0\})$ is a homothetically embedded $(\mathbb{S}^{n-1}/\Gamma)/\langle\sigma\rangle$ in $(\mathcal{O},g(t))$. Fix any $t$ and write $g=g(t)$. From Proposition A.4 in [27] we have the following
@@
(cf. Proposition 6.17 in [11] and Proposition 2.6 in [27]) Let $n\geq 4$ and $\varepsilon_{4}=\varepsilon_{4}(n)$ be as in Proposition 2.5. Let $(\mathcal{O},g)$ be a complete, noncompact orbifold diffeomorphic to $\mathbb{R}^{n}//\Gamma$ with strictly PIC and weakly PIC2 everywhere and strictly PIC2 at a point. Assume that $N$ is a $2\varepsilon_{4}$-neck in $\mathcal{O}$ diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$, and $\Sigma$ is a central cross section of $N$. When $\Gamma$ is nontrivial we assume that $\Sigma$ bounds a compact domain $\Omega$ in $\mathcal{O}$ whose interior contains the unique singular point of $\mathcal{O}$. Then if $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ is a $2\varepsilon_{4}$-homothety coming from the $2\varepsilon_{4}$-neck structure of $N$, $f$ extends to a diffeomorphism $F:D^{n}//\Gamma\rightarrow\Omega$.
@@
Definition. (cf. [11] and [27]) Let $\varepsilon_{0}=\varepsilon_{0}(n)$ be a small positive constant and $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$. Let $(\mathcal{O},g)$ be an $n$-dimensional Riemannian orbifold with at most isolated singularities. Given a point $x_{0}\in\mathcal{O}$, an open subset $U$ of $\mathcal{O}$ is an $\varepsilon$-cap centered at $x_{0}$ if $U$ is a topological cap and $U\setminus V$ is an $\varepsilon$-neck, where $V$ is a compact domain in $U$ diffeomorphic to some $D^{n}//\Gamma$ or $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B$ or $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B$ with $\partial V$ a central cross section of an $\varepsilon$-neck contained in $U$ (i.e., $\partial V$ is the image of $\mathbb{S}^{n-1}/\Gamma\times\{0\}$ under the diffeomorphism associated to the $\varepsilon$-neck) and with $x_{0}\in\text{Int}\hskip 1.42262ptV$, (we will call $V$ a core of the cap $U$,) and in addition,
1. when $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$, if $\Sigma$ is a central cross section of an $\varepsilon_{0}$-neck in $U\setminus V$ with an $\varepsilon_{0}$-homothety $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ coming from the $\varepsilon_{0}$-neck structure, $\Sigma$ bounds a compact domain $\Omega$ in $U$, and $f$ extends to a diffeomorphism $F:D^{n}//\Gamma\rightarrow\Omega$;
@@
(cf. Theorem 6.18 in [11], Theorem 3.9 in [15], and Proposition 2.7 in [27]) Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$ and $C_{2}=C_{2}(n,\theta,\varepsilon)$, such that given any noncompact orbifold ancient $\kappa$-solution $(\mathcal{O},g(t))$ of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}\text{id}\in C_{\text{PIC}}$ and is not locally isometric to an evolving shrinking round cylinder, for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$, which falls into one of the following categories:
(a) $U$ is a $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x_{0},t_{0})$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ with any cross-section of the $\varepsilon$-neck contained in the end of it diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$
for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$; moreover, the scalar curvature in $U$ at time $t_{0}$ is (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, and
$vol_{g(t_{0})}(U)>(C_{2}|\Gamma|)^{-1}R(x_{0},t_{0})^{-n/2}.$